OpenStudy (anonymous):

find the exact value : sin (-620 degrees)

OpenStudy (raden):

are u sure that is -620 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I already know it equals into 100 degrees, what I don't understand it actually finding the exact value for sin100 degrees.

OpenStudy (raden):

yeah, 100 is not the special degrees.. need to long time to get its exact

OpenStudy (anonymous):

I don't quite understand what your saying, I'm talking about Sum and Difference Identities.

OpenStudy (raden):

i think u want simplify it by using Sum or Difference Identities, right ? sin(-620) = -sin(620) = -sin(2*360-100) = -sin(-100) = (-)(-)sin(100) = sin100 if u want simplify again, it can be sin100 = sin(90+10) = -cos10

OpenStudy (anonymous):

It's +100, not -100

OpenStudy (raden):

10 degree is not the special angle too, in trigono we can getting its exact value, but too longgggg :)

OpenStudy (anonymous):

\(-620^\circ + 2\cdot 360^\circ = 100^\circ \)

OpenStudy (raden):

i used identity sin(-x) = -sinx, first @wio

OpenStudy (anonymous):

Sorry I think I wasn't clear about this question at the start. I have to use the formula: "sin(a +B) = sina cosB + cosa sinB " to find the exact value of 100 degrees

OpenStudy (anonymous):

I can't use values such as 30, 45, or 60 degrees because they don't equal to 100 degrees

OpenStudy (anonymous):

When you try to add 2 of them to eachother

OpenStudy (raden):

yes, we cant use 30,45,60 deg! better searching google :)

OpenStudy (anonymous):

I did! Which resulted hurting my hand from smacking it against my desk...

OpenStudy (raden):

maybe, this file useful for u

OpenStudy (raden):

and if ur question is -630 not -620, we can get it sin(-630) = -sin(630) = -sin(360+270) = -sin(270) = -(-1) = 1